# Newcastle-upon-Tyne, UK

### v2

@@ 1 Introduction
Modern mobile devices come equipped with an array of embedded sensors—accelerometers, gyroscopes, magnetometers, and others—that capture continuous motion and environmental data at fine temporal granularity. This rich sensor data has enabled applications from activity recognition to context-aware computing. More recently, research has proposed leveraging these signals for security-critical tasks such as cryptographic key generation, zero-interaction device pairing, and continuous authentication. A crucial yet under-explored assumption underpins such designs: sensor data provides sufficient unpredictability to thwart adversarial inference. Traditional “shake-to-pair” protocols [1] rely on motion patterns to establish secure communication between co-located devices, while newer techniques incorporate ambient phenomena—characteristics of magnetic fields [2] and thermal fluctuations [3]—to mitigate relay attacks or reduce user-driven authentication prompts. Despite these advancements, fundamental questions remain. Multi-modal sensing is often advocated to counter sensor-specific weaknesses [2, 4, 5, 6], but the quantitative security benefits of combining multiple sensors have not been rigorously evaluated. Many existing studies rely on heuristic assessments or machine learning classifiers, e.g. [4, 6, 7, 2, 5], that do not address core cryptographic requirements: how much entropy do sensors truly provide, and whether multi-modal combinations yield proportional security gains or suffer from inter-sensor correlations.
@@ 2.1 Mobile Sensor-based Security Applications
The notion of ‘hardness’ is typically inferred based on a machine learning-based analysis; it is not grounded in standard quantitative methods used in the field of authentication, such as Shannon [18] and min-entropy [19]. A general approach to follows one of those given by Shrestha et al. [20]. Here, sensor data is collected from $N$ users from which various features are extracted in the time or frequency domain (e.g. cross-correlation, spectral energy, Hamming, Euclidean and mean-absolute distances) [4, 6]. Calculating false positive (FPR) and negative (FNR) rates [1], precision and recall [12], EERs [7, 6], Receiver-Operator Curves (ROC) [16], and accuracy [16] have been used to evaluate the effectiveness of distinguishing between legitimate and illegitimate samples. User studies have also been employed to evaluate the effectiveness of sensor-based authentication systems [11, 1]. However, machine learning-based evaluation metrics do not measure intrinsic entropy—a crucial factor in determining whether sensor data is a suitable source of randomness for security applications. Some work has attempted to assess the entropy of sensor signals, but this represents a minority. T2Pair by Li et al. [21] 32.3–38.5 bits of Shannon entropy, with a refinement by Wu et al. [22] reporting 51–54 bits of Shannon entropy. Even in the best cases, this is low relative to modern cryptographic standards.
@@ 3.2 High-level Methodology
We aim to determine the global (i.e., population-level) entropy characteristics of various mobile sensors under ordinary usage conditions, rather than focusing on per-user or scenario-specific differences. This choice reflects common real-world deployments, which often must accommodate a wide range of behaviors and environments with minimal tailoring. Our approach involves five main stages:
@@ 4.3 Multi-modal Sensors
Our entropy analysis framework systematically evaluates the joint entropy over all valid combinations of sensor modalities. The powerset of the sensor set is generated (with restrictions on minimum and maximum combination lengths) and processed in parallel using Python’s multiprocessing module. Each subset is independently analysed: quantising the data, inferring a Chow–Liu tree, and computing the suite of entropy metrics. Processing all four datasets took approximately 22 hours on our workstation with an Intel i7-6700K (8M cache, 4.20 GHz) and 32 GB RAM on Ubuntu 24.04.

### v3

@@ 1 Introduction
Modern mobile devices come equipped with an array of embedded sensors—accelerometers, gyroscopes, magnetometers, and others—that capture continuous motion and environmental data at fine temporal granularity. This rich sensor data has enabled applications from activity recognition to context-aware computing. More recently, research has proposed leveraging these signals for security-critical tasks such as cryptographic key generation, zero-interaction device pairing, and continuous authentication. A crucial yet under-explored assumption underpins such designs: sensor data provides sufficient unpredictability to thwart adversarial inference. Traditional “shake-to-pair” protocols [1] rely on motion patterns to establish secure communication between co-located devices, while other methods have incorporated ambient phenomena, such as characteristics of magnetic fields and thermal fluctuations, to mitigate relay attacks [2, 3, 4, 5, 6] or reduce user-driven authentication prompts [7, 8, 9, 10, 11]. Despite these advancements, fundamental questions remain. Multi-modal sensing is often advocated to counter sensor-specific weaknesses [2, 4, 12, 13], but the quantitative security benefits of combining multiple sensors have not been rigorously evaluated. Many existing studies rely on heuristic assessments or machine learning classifiers, e.g. [4, 13, 14, 2, 12], that do not address core security requirements: how much entropy do sensors truly provide, and whether multi-modal combinations yield proportional security gains or suffer from inter-sensor correlations.
@@ 2.1 Mobile Sensor-based Security Applications
The notion of ‘hardness’ is typically inferred based on a machine learning-based analysis; it is not grounded in standard quantitative methods used in the field of authentication, such as Shannon [21] and min-entropy [22]. A general approach to follows one of those given by Shrestha et al. [3]. Here, sensor data is collected from $N$ users from which various features are extracted in the time or frequency domain (e.g. cross-correlation, spectral energy, Hamming, Euclidean and mean-absolute distances) [4, 13]. Calculating false positive (FPR) and negative (FNR) rates [1], precision and recall [8], EERs [14, 13], Receiver-Operator Curves (ROC) [11], and accuracy [11] have been used to evaluate the effectiveness of distinguishing between legitimate and illegitimate samples. (User studies have also been employed to evaluate the effectiveness of sensor-based authentication systems [18, 1].) We observe that machine learning-based evaluation metrics do not directly measure entropy, which is a crucial factor in determining whether sensor data is a suitable source of randomness for security applications. Some work has attempted to assess the entropy of sensor signals, but this represents a minority. T2Pair by Li et al. [23] 32.3–38.5 bits of Shannon entropy, with a refinement by Wu et al. [24] reporting 51–54 bits of Shannon entropy. Even in the best cases, this is low relative to modern cryptographic standards.
@@ 3.2 High-level Methodology
We aim to determine the global (i.e., population-level) entropy characteristics of various mobile sensors under ordinary usage conditions, rather than focusing on per-user or scenario-specific differences. This choice reflects common real-world deployments, which must accommodate a wide range of behaviors and environments. Our approach involves five main stages:
@@ 4.3 Multi-modal Sensors
Our entropy analysis framework systematically evaluates the joint entropy over all valid combinations of sensor modalities. The powerset of the sensor set is generated and processed in parallel using Python’s multiprocessing module. Each subset is independently analysed: quantising the data, inferring a Chow–Liu tree, and computing the suite of entropy metrics. Processing all four datasets took approximately 22 hours on our workstation with an Intel i7-6700K (8M cache, 4.20 GHz) and 32 GB RAM on Ubuntu 24.04.